OpenStudy (anonymous):

Find the value of 2√(3/25) A.1√(1/2) B.2√(3)/5<<< My Answer C.2√(3)/25 D.6/25 Is this right?

OpenStudy (anonymous):

is the entire fraction under the radical?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

B. is right

OpenStudy (anonymous):

I have five more questions that I wanted to get checked, should I post them here, or individually?

OpenStudy (anonymous):

whatever you want

OpenStudy (anonymous):

okay :) here's the next one: Find the value of √(8)/3√(4) A.6 B.2/3<<< My Answer C.√(2)/3 D.3√(2) for this one, I thought it was C at first, but for some reason I think its B?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

okay, so it was C, haha :) here's the next one: Find the value of √(24)/4√(6) in simplest form A.1 B.1/2<<< My Answer C.√(4)/4 D.4√(4)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

:) yay! here's the last one im gonna post, the others I have, im positive there right, so not needed.. Find the value of 7x^2/√(x^5) A.7√(x)/x<<< My Answer B.7/x C.7√(x) D.7x/√(x)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

are u sure, this one did confuse me a little, I solved it multiple times and I got A both times, but that's why I posted it so thanks

OpenStudy (anonymous):

i am sure

OpenStudy (anonymous):

okay :)

jhonyy9 (jhonyy9):

this will be 7/sqrtx - this posibillity not is there

OpenStudy (anonymous):

see c

OpenStudy (anonymous):

oh you are right

OpenStudy (anonymous):

renee hold on . I misread c

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

renee you are right it is a. johny is right but when you take his answer and get the rad out of the denominator you get answer a

OpenStudy (anonymous):

:) yay! I thought so... thanks :D

jhonyy9 (jhonyy9):

do you understand why ?

jhonyy9 (jhonyy9):

sqrtx^5 =sqrt(x^4 *x) =x^2sqrtx

OpenStudy (anonymous):

yes! thanks again to the both of you for your help!